calculate the angular frequency of hinged rigid body.
1
Expert's answer
2020-09-17T14:30:05-0400

Oscillations of hinged rigid body are described by the model of physical pendulum.

For mathematical pendulum "\displaystyle T = 2 \pi \sqrt{\frac{L}{g}}" where L is a length of pendulum.

For physical pendulum we can introduce equivalent length L - the distance from the pivot to the centre of oscillation.

"\displaystyle L = \frac{I}{mR}"

where where I is the moment of inertia of the pendulum about the pivot point, m is the mass of the pendulum, and R is the distance between the pivot point and the centre of the mass.

So, "\displaystyle T = 2\pi \sqrt{\frac{I}{mgR}}" .

Angular frequency "\displaystyle \omega = \frac{2\pi}{T} =\sqrt{ \frac{mgR}{I}}" .